Please help m solve the following:
((x-5)/(x+3))= ((2(x+1))/(x+3))
Thank you

Answer by Boreal(15235)

((x-5)/(x+3))= ((2(x+1))/(x+3))
The denominators cancel each other
x-5=2(x+1)=2x+2
x=-7
check
-12/-4 (or 3) = 2(-6)/-4 (or 3)